# Glimpse: Generalized Local Imaging with MLPs

### v2

@@ III-B Adaptive Filtering for Noisy Measurements
The Ram–Lak high-pass filter is the optimal filter ${\mathbf{h}}$ for the FBP reconstruction in the case of complete noise-free measurements; see Appendix -C for a standard demonstration. In real applications, however, we always encounter noisy projections from a subset of angles. The Ram–Lak filter is then suboptimal and typically degrades the reconstruction quality as it amplifies high-frequency noise. Alternative filters with lower amplitudes in high frequencies like Shepp–Logan, cosine, and Hamming have been used to mitigate the noisy measurements, but they are all ad hoc choices. It is advantageous to adapt ${\mathbf{h}}$ to the specifics of noise and sampling strategy in the target application. To design this task-specific filter, we let $\text{MLP}_{\mathrm{\bm{\theta}}}$ take as input the filtered sinogram $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$ and consider the filter ${\mathbf{h}}$ (in Fourier space) as trainable parameters to be optimized during training. This allows us to automatically learn a noise-adaptive filter from data, again with almost no additional computational cost.
III-C Resolution-agnostic Memory Usage in Training
@@
To simplify notation, we denote the entire Glimpse pipeline described above by $\hat{{\mathbf{f}}}({\mathbf{x}})=\text{{Glimpse}}_{\phi}({\mathbf{x}},{\mathbf{s}})$. The inputs are the target pixel coordinates ${\mathbf{x}}=(x,y)$ and the sinogram ${\mathbf{s}}$; the output is an estimate of ${\mathbf{f}}(x,y)$. The parameters $\phi$ denote the trainable parameters of Glimpse including the MLP weights $\theta$, the projection angles $\{\alpha_{m}\}_{m=1}^{r}$ (see Appendix -B), the adaptive filter ${\mathbf{h}}$ and the window receptive field scale $d$. We consider a set of training data $\{({\mathbf{s}}_{i},{\mathbf{f}}_{i})\}_{i=1}^{L}$ from the noisy sinograms and images. We optimize the Glimpse parameters $\phi$ using gradient-based optimization to minimize
@@
At inference time, we simply evaluate the image intensity at any pixel as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{{Glimpse}}_{\phi^{*}}({\mathbf{x}},{\mathbf{s}}_{\text{test}})$. One major advantage of Glimpse compared to CNNs like U-Net and LPD is its memory and compute complexity. CNN-based models exhibit memory requirements that scale directly with image resolution, making them prohibitively expensive for realistic image resolutions. As shown in (10), Glimpse can be trained using stochastic gradient-based optimizers with the flexibility to select mini-batches from both the objects and pixels. This adaptability in mini-batch pixel selection grants a memory footprint agnostic to resolution making Glimpse suitable for training on realistic image resolutions like $1024\times 1024$ and higher.
IV Experiments
We simulate parallel-beam X-ray CT with $r=30$ projections uniformly distributed around the object with additive Gaussian noise to reach a signal-to-noise ratio (SNR) of 30 dB. The reconstruction quality is quantified using the peak signal-to-noise ratio (PSNR) and Structural Similarity Index (SSIM) [40]. We compare the performance of Glimpse with successful CNN-based models: U-Net [2], learned gradient scheme (LGS) [7] and learned primal-dual (LPD) [8] for sparse view CT image reconstruction. We use 35820 training samples of chest images from the LoDoPaB-CT dataset [41] in resolution $128\times 128$. Model performance is assessed on 64 in-distribution test samples of chest images, while 16 OOD brain images [42] are included to evaluate the generalization capability of the models. For further information regarding the network architectures and training details please refer to Section -A.
@@ IV Experiments
In Section IV-A, we compare Glimpse to CNN-based models for sparse view CT image reconstruction on both in-distribution and OOD data. In Section IV-B, we analyze the computational efficiency of the aforementioned models. We analyze the learned filters ${\mathbf{h}}$ across different measurement noise levels in Section IV-C. In Appendix -B we consider the uncalibrated and blind scenarios.
IV-A Sparse view CT Image Reconstruction
Figure 4: Performance of different models trained on training data of chest images evaluated on in-distribution and OOD samples for sparse view CT image reconstruction. Glimpse has excellent performance on OOD data due to its localized MLP, significantly better than U-Net [2] and LGS [7] and comparable with LPD [8].
TABLE I: Comparison of different models for sparse view CT image reconstruction
(a) The reconstruction quality averaged on 64 test samples
In-distribution (chest) Out-of-distribution (brain)
PSNR SSIM PSNR SSIM
FBP [14] 17.0 0.17 17.1 0.22
U-Net [2] 30.1 0.84 15.1 0.28
LGS [7] 30.9 0.84 20.5 0.54
LPD [8] 31.6 0.86 25.5 0.76
Glimpse 30.9 0.84 25.1 0.79
(b) Memory usage and training time (batch size 64)
Glimpse U-Net [2] LGS [7] LPD [8]
Num params 900k 7800k 19k 400k
$128\times 128$ 4GB / 114s 6GB / 34s 4GB / 384s 13GB / 963s
$256\times 256$ 4GB / 123s 16GB / 117s 13GB / 575s 41GB / 1517s
$512\times 512$ 4GB / 185s 53GB / 460s 45GB / 1682s $>80$GB
$1024\times 1024$ 5GB / 419s $>80$GB $>80$GB $>80$GB
The upper row of Figure 4 and Table I(a) show the performance of different models on in-distribution test samples of chest images. This experiment shows that Glimpse, by leveraging only a single MLP network, can outperform successful CNNs like U-Net and achieve comparable performance with LGS and LPD methods.
The lower row of Figure 4 and Table I(a) shows a comparison of the performance of various models trained on chest images when applied to OOD brain images. This experiment demonstrates that while U-Net excels on in-distribution samples, its performance significantly deteriorates on OOD data.
On the contrary, Glimpse shows strong performance on OOD data. Although LPD’s performance on OOD data is sometimes comparable or slightly better than that of Glimpse, it comes at a very high memory and compute cost due to the repeated application of the forward operator and its adjoint in the network architecture; we analyze this in the next section.
IV-B Computational Efficiency
The fact that LPD far outperforms U-Net on OOD data is a testament to the benefits of incorporating the forward operator in the architecture. On the other hand, as evident from Table I(b), it comes at the cost of unfavorable training time and memory footprint which rapidly worsens with resolution. Table I(b) shows that CNN-based models may become impractical already at resolutions like $512\times 512$, even on GPUs with 80GB memory.
On the other hand, Glimpse is computationally efficient; the memory usage remains almost unaffected by image resolution. Remarkably, Glimpse can be trained with only 5GB memory in less than a day, even when dealing with resolutions of $1024\times 1024$ and higher. Figure 5 shows the performance of Glimpse on in-distribution and OOD samples in resolution $512\times 512$ where we considered 40dB measurement noise. This experiment demonstrates that a relatively small MLP, with almost 10 times fewer parameters than a standard U-Net, can achieve strong performance in realistic high resolutions while maintaining a rather modest memory footprint.
@@ IV-B Computational Efficiency
Figure 5: Glimpse’s performance in resolution $512\times 512$ trained on chest training data with $r=30$ projections and 40dB noise; Glimpse requires only 4GB memory and can be trained in less than 10 hours on a single GPU.
IV-C Learned Filter
Figure 6: The learned filter for datasets with different noise levels, all the filtered are initialized by Ram-Lak filter in Glimpse architecture. By increasing the noise level, the filter assigns smaller amplitudes for high-frequencies to suppress the noise and aligns with the optimality of the Ram-Lak filter for noise-free complete measurements shown in Section -C.
In this section, we analyze the behavior of the learned filter obtained through training of Glimpse introduced in section III-B across datasets with different measurement noise levels. This analysis provides useful signal processing insights into how the properties of the learned filter are influenced by varying noise levels.
In Figure 6 we show the frequency response of the learned filters, alongside with standard hand-crafted filters such as Ram-Lak, Shepp-Logan, and Hamming filters. These learned filters are derived from Glimpse training on datasets characterized by different levels of noise.
As expected from the discussion in Appendix -C, the learned filter for noise-free measurements is similar to the Ram–Lak filter, with a relatively high amplitude in high frequencies. As the noise level increases (by decreasing the noise SNR), the filter progressively takes smaller values in high frequencies to suppress the noise. This confirms that Glimpse can autonomously adapt the characteristics of the filter according to the noise level observed in the training data.
V Limitations and Conclusion
We used a natural notion of locality for CT which is adapted to sinogram geometry. This is different from CNNs that reconstruct the image as a whole, and where the notion of locality (at small scales) is in the sense of real image space. Our approach adopts a coordinate-based strategy, focusing on processing the sinusoidally-shaped regions of the sinogram associated with pixels using a small MLP. Our results demonstrate that this localized processing framework does significantly improve robustness to OOD data while maintaining nearly constant memory requirements across different resolutions, being computationally efficient even at realistic high resolutions. The differentiable architecture makes it easy to combine Glimpse with other deep learning pipelines. Concretely, we demonstrated the possibility to learn sensor geometry for uncalibrated systems and to adapt the learned filter according to the noise level and projection sparsity.
While the memory required by Glimpse varies little with image resolution, a drawback of the current scheme is that memory and computing costs increase as the number of projections $r$. A possible alternative to the standard MLP architecture which is the culprit for this is to use mixture-of-experts layers [43, 44, 45], which selectively employ smaller MLPs for processing inputs. This mixture-of-experts approach is an effective drop-in replacement for standard MLP layers of language transformers [46] and vision transformers [24].
Glimpse could be integrated with various imaging problems involving line integrals in forward operators such as photoacoustic [47, 48] and cryo-electron tomography (cryoET) [49, 50]. Its future full-3D adaptation may yield efficient architectures which resolve the fundamental memory issues with applications of deep learning in 3D medical imaging. This extension is particularly interesting given the ability of Glimpse to operate locally and its near-fixed memory requirement across resolution, which makes it an ideal choice for large 3D objects.
-A Network Architecture and Training Details
@@ V Limitations and Conclusion
For Glimpse architecture, we use an MLP network comprising 9 hidden layers, each with dimensions [256, 256, 256, 256, 128, 128, 128, 64, 64] with ReLu activations. The input to the MLP network consists of sinusoidal curves sampled from $K=9^{2}$ neighboring pixels. To prevent edge artifact of the circular convolution, we apply zero-padding with a size of 512 to the sinogram before applying the filter ${\mathbf{h}}$. Linear interpolation is used in (7). For the experiment in resolution $512\times 512$ in Section IV-B, we use a larger network with hidden layer dimensions [1024, 1024, 1024, 512, 512, 512, 256, 256] to enhance the quality of reconstructions.
We implement our model in PyTorch [51] on a machine equipped with a Nvidia A100 GPU with 80GB of memory to train the different architectures. We report the maximum capacity of the graphics card during training and the time needed to complete the training. All models in Section IV were trained for 200 epochs with MSE loss using the Adam optimizer [52]. A learning rate $10^{-4}$ was used for Glimpse and U-Net, while LGS and LPD were trained with a learning rate $10^{-3}$. All models were trained with batch size 64. In the case of Glimpse, for each mini-batch of random objects, we performed optimization on a random mini-batch of 512 pixels 3 times.
-B Learned Sensor Geometry

### v3

@@ 3.4 Adaptive Filtering for Noisy Measurements
The Ram-Lak high-pass filter is the optimal filter ${\mathbf{h}}$ for the FBP reconstruction in the case of complete noise-free measurements; see Appendix .4 for a standard demonstration. In real applications, however, we always encounter noisy projections from a subset of angles. The Ram-Lak filter is then suboptimal and typically degrades the reconstruction quality as it amplifies high-frequency noise. Alternative filters with lower amplitudes in high frequencies like Shepp-Logan, Cosine, and Hamming have been used to address this, but they are all ad hoc choices. It is advantageous to adapt ${\mathbf{h}}$ to the specifics of noise and sampling strategy in the target application. To design this task-specific filter, we consider the coefficients of the filter ${\mathbf{h}}$ (in the frequency domain) as trainable parameters to be optimized during training as depicted in Figure 1. This allows us to automatically learn a noise-adaptive filter from data, again with almost no additional computational cost.
3.5 Resolution-agnostic Memory Usage in Training
@@
Glimpse is fully differentiable which enables the optimization of the receptive field scale, filter parameters, and MLP weights via backpropagation during training. To simplify notation, we denote the entire described Glimpse pipeline by $\hat{{\mathbf{f}}}({\mathbf{x}})=\text{{Glimpse}}_{\phi}({\mathbf{x}},{\mathbf{s}})$. The inputs are the target pixel coordinates ${\mathbf{x}}=(x,y)$ and the sinogram ${\mathbf{s}}$; the output is an estimate of ${\mathbf{f}}(x,y)$. The parameters $\phi$ denote the trainable parameters of Glimpse including the MLP weights $\theta$, the projection angles $\{\alpha_{m}\}_{m=1}^{r}$ (see Appendix .3), the adaptive filter ${\mathbf{h}}$ and the window receptive field scale $d$. We consider a set of training data $\{({\mathbf{s}}_{i},{\mathbf{f}}_{i})\}_{i=1}^{L}$ from the noisy sinograms and images. We optimize the Glimpse parameters $\phi$ using gradient-based optimization by minimizing
@@
At inference time, we simply evaluate the image intensity at any pixel as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{{Glimpse}}_{\phi^{*}}({\mathbf{x}},{\mathbf{s}}_{\text{test}})$. One major advantage of Glimpse compared to CNNs like U-Net and LPD is its low memory and compute complexity. Memory requirements of CNN-based models scale steeply with image resolution, making them prohibitively expensive for realistic resolutions. As shown in (10), Glimpse can be trained using stochastic gradient-based optimizers with the flexibility to select mini-batches from both the objects and pixels thanks to its coordinate-based design. This leads to a memory footprint nearly agnostic to resolution, which makes Glimpse suitable for training on realistic image resolutions of $1024\times 1024$ and higher.
4 Experiments
We benchmark Glimpse against successful CNN-based baselines for sparse-view CT reconstruction: U-Net [2], iRadonMAP [29] with U-Net as the post-processing CNN, learned gradient scheme (LGS) [9] and learned primal-dual (LPD) [10]. For a thorough comparison we created two additional baselines: 1) iRadonMAP-ff: in the original iRadonMAP, the filter ${\mathbf{h}}$ in (6) is replaced with an MLP architecture. Here, we consider iRadonMAP-ff which rather uses the learnable Fourier filter ${\mathbf{h}}$ introduced in Section 3.4, allowing us to ablate the effects of different filtering procedures; 2) iRadonMAP-ffnu: the original iRadonMAP employs a post-processing CNN to enhance reconstruction quality. To assess the performance of the linear model alone, we consider iRadonMAP-ffnu, which excludes the CNN. This comparison with Glimpse helps us understand the significance of our non-linear mapping $\text{NN}_{\theta}$ and the inclusion of neighboring pixels. The reconstruction quality is quantified using the peak signal-to-noise ratio (PSNR) and Structural Similarity Index (SSIM) [60]. Bottom left windows in Figures show the PSNR between the reconstructed image and the ground truth.
We implement all models in PyTorch [61] on a machine equipped with a Nvidia A100 GPU with 80GB memory. All models were trained for 200 epochs with MSE loss using the Adam optimizer [62]. A learning rate of $10^{-4}$ was used for Glimpse, U-Net and iRadonMAP, and of $10^{-3}$ for LGS and LPD. All models were trained with batch size 64. For Glimpse, for each mini-batch of random targets, we ran optimization on a random mini-batch of 512 pixels 3 times.
@@ 4 Experiments
In Section 4.1, we compare Glimpse to CNN-based models for sparse-view CT reconstruction on both in-distribution and OOD data. In Section 4.2, we analyze the computational efficiency of the aforementioned models. We analyze the learned filters ${\mathbf{h}}$ across different measurement noise levels in Section 4.3. We study the influence of the number of projections and neighboring pixels in Sections 4.4 and 4.5. Finally, in Appendix .3, we present our method for learning the projection angles jointly with the image reconstruction to address uncalibrated and blind scenarios.
4.1 Sparse view CT Image Reconstruction
@@
Figure 6: Performance of different models trained on training data of chest images and evaluated on in-distribution and OOD samples. Glimpse shows very strong performance on OOD data, significantly better than U-Net [2], iRadonMAP [29], LGS [9] and comparable with LPD [10]. We indicate PSNRs between the reconstructions and the ground truth.
Table 1: Comparison of different models for sparse view CT. The reconstruction quality is calculated on 64 test samples.
Methods Num params In-distribution (chest) Out-of-distribution (brain)
PSNR SSIM PSNR SSIM
FBP [16] 00 $17.0\pm 1.9$ $0.17\pm 0.06$ $17.1\pm 1.3$ $0.22\pm 0.02$
U-Net [2] $7800$k $30.1\pm 1.4$ $0.84\pm 0.02$ $15.1\pm 1.8$ $0.28\pm 0.03$
iRadonMAP [29] $8400$k $28.5\pm 1.3$ $0.80\pm 0.03$ $13.4\pm 2.1$ $0.23\pm 0.06$
iRadonMAP-ff $8200$k $30.1\pm 1.3$ $0.83\pm 0.02$ $14.2\pm 1.6$ $0.25\pm 0.04$
iRadonMAP-ffnu $500$k $25.2\pm 1.5$ $0.64\pm 0.03$ $19.5\pm 1.9$ $0.36\pm 0.07$
LGS [9] $19$k $30.9\pm 1.4$ $0.84\pm 0.02$ $20.5\pm 7.7$ $0.54\pm 0.31$
LPD [10] $400$k $\textbf{31.6}\pm 1.4$ $\textbf{0.86}\pm 0.02$ $\textbf{25.5}\pm 2.6$ $0.76\pm 0.06$
Glimpse (MLP) $900$k $30.9\pm 1.4$ $0.84\pm 0.02$ $25.1\pm 2.3$ $\textbf{0.79}\pm 0.05$
Glimpse (MultiMLP) $900$k $31.0\pm 1.4$ $0.84\pm 0.02$ $25.0\pm 2.3$ $0.77\pm 0.05$
We simulate parallel-beam X-ray CT with $r=30$ projections uniformly distributed around the object with additive Gaussian noise to reach a signal-to-noise ratio (SNR) of 30 dB. Model performance is assessed on 64 in-distribution test samples of chest images, while 16 OOD brain images [63] are included to evaluate the generalization capability of the models.
@@
Glimpse (MLP) uses an MLP with 9 hidden layers of dimensions [256, 256, 256, 256, 128, 128, 128, 64, 64], with ReLU activations. Glimpse (MultiMLP) consists of nine small MLP blocks, each with three hidden layers of size 128. The outputs of these MLPs are then combined using an additional MLP with the same architecture. To ensure a fair comparison, both Glimpse (MLP) and Glimpse (MultiMLP) are designed to have a comparable number of trainable parameters. The input to the MLP network consists of sinusoidal curves sampled from $K=9^{2}$ neighboring pixels. To prevent boundary cross talk due to circular convolution (since we implement an unconstrained discrete Fourier transform multiplier), we apply zero-padding with a size of 512 to the sinogram before applying the filter ${\mathbf{h}}$. Linear interpolation is used in (7).
4.1.1 Training data of chest images
We use 35820 training samples of chest images from the LoDoPaB-CT dataset [64] in resolution $128\times 128$. Figure 6(a) and Table 1 show the performance of different models on in-distribution test samples of chest images. We see that Glimpse (MLP) and Glimpse (MultiMLP) outperform successful CNNs like U-Net and iRadonMAP and achieve comparable performance with LGS and LPD methods, all while using simple MLPs.
Figure 6(b) and Table 1 compare the various models trained on chest images and applied to OOD brain images. This experiment demonstrates that while U-Net, iRadonMAP and iRadonMAP-ff excel on in-distribution samples, their performance significantly deteriorates on OOD data.
By contrast, Glimpse (MLP) shows strong performance on OOD data. Glimpse (MultiMLP) achieves comparable performance with Glimpse (MLP) which showcases the suitability of the new MultiMLP architecture. Although LPD’s performance on OOD data is sometimes comparable or slightly better than that of Glimpse, it comes at an extremely high memory and compute cost; we analyze this further in Section 4.2.
Table 1 also highlights the superior performance of Glimpse compared to iRadonMAP and its variants, particularly iRadonMAP-ffnu, which excludes the post-processing CNN. This can be explained by two key factors: (1) Unlike iRadonMAP, which extracts a single sinusoidal curve per pixel, Glimpse also processes neighboring pixels, enabling significantly better reconstructions; and (2) while iRadonMAP-ffnu uses a linear transformation for local neighborhood processing, Glimpse leverages a much more expressive non-linear mapping via MLPs.
On the other hand, iRadonMAP and iRadonMAP-ff show better reconstruction on in-distribution chest data but generalize poorly compared to the local processing iRadonMAP-ffnu. This is due to the post-processing CNN in iRadonMAP and iRadonMAP-ff, which negatively impacts generalization. Finally, the filter in iRadonMAP-ff outperforms the MLP filter in the original version, demonstrating the advantage of simple linear filtering, as discussed in Section 3.4.
Table 2: Comparison of different models for sparse view CT image reconstruction; the reconstruction quality is calculated on 64 test samples.
Datasets Num samples Chest Brain
PSNR SSIM PSNR SSIM
Chest [64] $35820$ 30.9 0.84 $25.1$ $0.79$
DIV2K [65] $800$ $27.8$ $0.75$ $23.3$ $0.65$
CelebA-HQ [66] $30000$ $28.8$ $0.79$ 25.3 0.80
4.1.2 Training data of natural images
The robustness of Glimpse to distribution shift motivates an experiment to examine the impact of the training dataset on performance. For this purpose, we consider two distinct datasets of natural images: (1) DIV2K [65], with 800 high-quality natural images, and (2) CelebA-HQ [66], with 30,000 high-resolution images of human faces. Except the training dataset, the network architecture and the training details are the same as Section 4.1.1. Table 2 presents the performance of Glimpse trained on these datasets and applied to chest and brain medical images. Notably, CelebA-HQ, despite being visually unrelated to medical images, trains Glimpse as effectively as the chest dataset. By contrast, training with a smaller dataset like DIV2K results in a significant drop in reconstruction quality, highlighting the importance of large high-quality data for improving model generalization.
4.2 Computational Efficiency
(a) Memory footprint (10 images)
(b) Inference time (10 images)
Figure 7: The memory and time requirements during inference for different models.
The fact that LPD far outperforms U-Net on OOD data is a testament to the benefits of incorporating the forward operator in the architecture. However, evaluating the Radon transform and its adjoint can become prohibitively expensive for large images, as it implies storing multiple copies of the same size as the original image. It can be partially mitigated by reducing the number of iterations in the associated iterative reconstruction scheme but at the cost of a significant deterioration in reconstruction quality. In this section, we compare the training memory and time requirements of different models at different resolutions, for 500 iterations with batch size 64. We report the maximum use of GPU memory and the time needed to complete the training and inference. As evident from Figure 2, the success of LPD and LGS comes at the cost of very unfavorable training memory and time complexity which rapidly worsens with resolution. On the other hand, the memory needed to train Glimpse is almost independent from image resolution. Remarkably, Glimpse needs only 5GB memory to train on $1024\times 1024$ images—less than 1/16 of the memory typically needed by standard CNNs for CT image reconstruction. This makes Glimpse suitable for high dimensional reconstruction tasks in real-world applications.
We next compare the computational efficiency of various models during inference. With Glimpse, there is a trade-off between inference speed and memory usage: smaller batch sizes reduce memory consumption but slow down inference, whereas larger batch sizes enable faster inference at the cost of higher memory usage. In this experiment, we set the pixel batch size to 1024. Figure 7 presents the memory footprints and runtimes of different models for reconstructing 10 samples. Although Glimpse performs pixel-wise image synthesis, it remains comparable to other CNNs that recover the whole image at once. For further discussion on the computational cost and potential remedies, please refer to Section 5.1.
Finally, we study the performance of Glimpse (MultiMLP) on higher-resolution CT reconstruction. We train on the LoDoPaB-CT dataset at resolution $512\times 512$, using 90 projections with 40dB measurement noise. For this experiment, we use a larger MultiMLP with hidden layer dimension 400 to enhance the quality of reconstructions. Figure 8 shows the performance of Glimpse on in-distribution and OOD samples, along with the pixel-wise absolute error maps between the reconstructions and ground truth images. This experiment demonstrates that our proposed framework can achieve strong performance in realistic high resolutions.
(a) In-distribution chest samples
(b) OOD brain samples
Figure 8: Glimpse’s performance in resolution $512\times 512$ trained on chest training data with $r=90$ projections and 40dB noise. We indicate PSNRs between the reconstructions and the ground truth along with the pixel-wise absolute error maps.
4.3 Learned Filter
Figure 9: The learned filter for datasets with different noise levels, all the filtered are initialized by Ram-Lak filter in Glimpse architecture. By increasing the noise level, the filter assigns smaller amplitudes for high-frequencies to suppress the noise and aligns with the optimality of the Ram-Lak filter for noise-free complete measurements shown in Section .4.
In this section, we study the learnable filter introduced in Section 3.4 across datasets with different measurement noise levels. This provides useful signal processing insights into how the properties of the learned filter are influenced by varying noise levels. In Figure 9 we show the frequency response of the learned filters, alongside standard hand-crafted filters such as Ram-Lak, Shepp-Logan, and Hamming. The learned filters are trained jointly with the MLPs in Glimpse. As expected (see also the discussion in Appendix .4), the learned filter for noise-free measurements is similar to the Ram-Lak filter, with a relatively high amplitude in high frequencies. As the noise level increases (by decreasing the noise SNR), the filter progressively takes smaller values in high frequencies to suppress the noise. This shows that Glimpse can indeed autonomously adapt the characteristics of the filter according to noise (and other characteristics) in the training data. We additionally observe that training Glimpse with a learnable filter leads to much faster convergence compared to a fixed filter (such as the Ram-Lak) while achieving comparable (or slightly better) reconstruction quality. Reconstructed images for different noise levels are presented in Figure 10.
(a) In-distribution chest samples
(b) Out-of-distribution brain samples
Figure 10: Glimpse performance on in-distribution and OOD data for different measurement noise levels with $r=30$ projections. We indicate PSNRs between the reconstructions and the ground truth.
4.4 Influence of the Number of Projections
As mentioned in Section 3.3, Glimpse (MultiMLP) can process measurements with large number of projections $r$. To show the effectivity of the proposed architecture, we study the performance of Glimpse (MultiMLP) for different number of projections while we have 30dB measurement noise. Separate Glimpse (MultiMLP) models were trained on datasets with varying numbers of projections. Figure 11 shows the reconstructions for different number of projections.
(a) In-distribution chest samples
(b) Out-of-distribution brain samples
Figure 11: Glimpse performance on in-distribution and OOD data for different number of projections with measurement noise 30dB. We indicate PSNRs between the reconstructions and the ground truth.
4.5 Influence of the Neighborhood Size
In this section, we analyze the significance of contextual information on Glimpse’s performance by varying the number of neighboring pixels (patch size) $K=C^{2}$. Table 3 presents the performance of Glimpse trained with different patch sizes $K$ on both in-distribution and out-of-distribution (OOD) samples. The results demonstrate that Glimpse with $K=3\times 3$ significantly outperforms the model without contextual information ($K=1$). Moreover, we see that the reconstruction quality tends to reach a saturation point beyond a certain patch size. This observation can inform the optimal choice of context size.
Table 3: Reconstruction quality in PSNR (dB) for Glimpse trained with various number of neighboring pixels.
Patch size ($K=C^{2}$) In-distribution OOD Num params
$1\times 1$ 25.6 18.3 280k
$3\times 3$ 30.2 24.2 345k
$5\times 5$ 30.7 24.9 470k
$7\times 7$ 30.8 24.9 650k
$9\times 9$ 30.9 25.1 900k
$11\times 1$ 11 30.9 25.0 1200k
5 Discussions and Conclusion
We have demonstrated that Glimpse — a neural network adapted to the geometry of computed tomography—can be much more robust, much more scalable, and much less data hungry CT reconstructions than the leading CNN-based (and model-based) methods. Our experiments substantiate the key claims made in the Introduction. First, by exploiting local sinusoidal patches in the sinogram, Glimpse handles out-of-distribution data more gracefully than leading CNN-based methods. Second, since training is done at the pixel level, Glimpse’s GPU memory usage remains nearly constant as the image resolution grows, making it scalable to 1024x1024 or higher without requiring prohibitively large hardware. Finally, the learnable filter and differentiable projection angles make Glimpse highly flexible in practice, able to handle noisy datasets and even uncalibrated systems where sensor geometry is only partially known. This last feat is facilitated by the robustness and numerical efficiency of Glimpse.
5.1 Limitations
Glimpse can be trained on GPUs with significantly smaller memory than baselines, which enables very high-dimensional image reconstruction, but its computational cost at inference scales with the number of pixels. Recent work [67, 68] has improved the efficiency of continuous image representation in INRs by increasing shared computations across coordinates, thereby reducing computational complexity. Adapting these methods within Glimpse could potentially decrease inference time. We note, however, that even with the current architecture inference is essentially real-time.
Another challenge is that memory and compute cost increase with the number of projections $r$. A possible alternative to the standard MLP or MultiMLP architectures which are the culprit for this is to use mixture-of-experts layers [69, 70, 71], which selectively employ smaller MLPs for processing inputs. This approach is an effective drop-in replacement for standard MLP layers of language transformers [72] and vision transformers [59]; we leave it to future work to test its effectiveness in local CT reconstruction.
Since the dimensionality of the MLP network is fixed, Glimpse can only process data with the specific number of projections it was trained on. This limitation is common in most deep-learning models for tomographic reconstruction, including model-based architectures like LPD and LGS. Here, however, it arises specifically from the MLP structure. Architectures such as transformers [72], which can process data sequentially, are likely the right solution.
5.2 Looking forward: locality for other imaging modalities
Glimpse can be generalized to various imaging problems where the forward operator involves line integrals, such as fan-beam computed tomography (CT) [51]. In fan-beam CT, X-rays diverge from a source point in a fan-shaped pattern as they pass through the object, a configuration commonly used in clinical CT scanners due to its efficiency in capturing larger areas. As detailed in [73, §5.11.6], although the fan-beam CT forward operator is more complex than that of parallel-beam CT, it retains a local structure that can be exploited to develop a local processing reconstruction pipeline, similar to Glimpse. Glimpse can also be extended to other imaging modalities with a local forward operator including photoacoustic [74, 75] and cryo-electron tomography (cryoET) [76, 77]. Its future full-3D adaptation may yield efficient architectures that resolve the fundamental memory issues with applications of deep learning in 3D medical imaging. This extension is particularly interesting given the ability of Glimpse to operate locally and its near-fixed memory requirement across resolution, which makes it a potentially strong choice for full 3D problems.
.3 Learned Sensor Geometry